# Behavior-Regularized Diffusion Policy Optimization for Offline Reinforcement Learning

### v1

@@ Appendix C Proof for the Main Results
The proof duplicates that in Haarnoja et al. (2018).
@@ Appendix D Continuous-Time Perspective of the Pathwise KL
Song et al. (2021) builds the connection between diffusion models and the reverse stochastic differential equations (SDEs) that gradually remove noises from data and transform a prior noise distribution into the target distribution. Based on the SDE understanding of diffusion models, we aim to investigate whether the introduced pathwise KL can be extended to continuous-time diffusion modeling.
We define the forward diffusion SDE, which perturbs data into prior distribution, as the following:
@@
$\operatorname{KL}\left[\mathbb{P^{\pi}}\|\mathbb{P}^{\nu}\right]$ $=\mathbb{E}_{\mathbb{P}^{\pi}}\left[\log\frac{\mathrm{d}\mathbb{P}^{\pi}}{\mathrm{d}\mathbb{P}^{\nu}}\right]$ (31)
$=\mathbb{E}_{\mathbb{P}^{\pi}}\left[\int_{0}^{T}\frac{[[f(x,t)-g(t)^{2}s^{\pi}(x,t)]-[f(x,t)-g(t)^{2}s^{\nu}(x,t)]]^{2}}{2g(t)^{2}}\mathrm{d}t\right]$
$=\mathbb{E}_{\mathbb{P}^{\pi}}\left[\int_{0}^{T}\frac{g(t)^{4}\|s^{\pi}(x,t)0s^{\nu}(x,t)\|^{2}}{2g(t)^{2}}\mathrm{d}t\right]$
@@
$\operatorname{KL}\left[\mathbb{P}^{\pi}\|\mathbb{P}^{\nu}\right]$ $=\mathbb{E}_{\mathbb{P}_{\pi}}\left[\int_{0}^{T}\frac{\beta(t)}{2}\|s^{\pi}(x,t)-s^{\nu}(x,t)\|^{2}\mathrm{d}t\right].$ (33)
Equation 31 presents the analytical form of the KL divergence between two reverse diffusion processes. In the following content, we continue to demonstrate that the pathwise KL introduced in this paper corresponds to Eq. (31) in the limit of $N\to\infty$.
@@
$\operatorname{KL}\left[p^{\pi}_{0:N}\|p^{\nu}_{0:N}\right]$ (34)
$=\mathbb{E}_{p^{\pi}_{0:N}}\left[\sum_{n=1}^{N}\operatorname{KL}\left[p^{\pi,x^{n}}_{n-1|n}(x^{n-1})\|p^{\nu,x^{n}}_{n-1|n}(x^{n-1})\right]\right]$
@@
$=\mathbb{E}_{p^{\pi}_{0:N}}\left[\sum_{n=1}^{N}\frac{\beta_{n}(1-\bar{\alpha}_{n})}{2(1-\bar{\alpha}_{n-1})(1-\beta_{n})}\|s_{n}^{\pi}(x)-s_{n}^{\nu}(x)\|^{2}\right].$
@@
$\lim_{N\to\infty}\operatorname{KL}\left[p^{\pi}_{0:N}\|p^{\nu}_{0:N}\right]$ $=\lim_{N\to\infty}\mathbb{E}_{p^{\pi}_{0:N}}\left[\sum_{n=1}^{N}\frac{\beta_{n}(1-\bar{\alpha}_{n})}{2(1-\bar{\alpha}_{n-1})(1-\beta_{n})}\|s_{n}^{\pi}(x)-s_{n}^{\nu}(x)\|^{2}\right]$ (35)
@@
$=\mathbb{E}_{p^{\pi}}\left[\int_{0}^{T}\frac{\beta(t)}{2}\|s^{\pi}(x(t),t)-s^{\nu}(x(t),t)\|^{2}\mathrm{d}t\right].$
@@ E.1 Discussions about Training Time
A systematic breakdown and comparison of the runtime between BDPO, DAC and Diffusion-QL is illustrated in Figure 10. We evaluate BDPO, DAC and Diffusion-QL with workstations equipped with NVIDIA RTX 4090 cards and the walker2d-medium-replay-v2 dataset.
@@
E.2 Generation Path On Synthetic 2D Tasks
We plot the generation results of BDPO on all synthetic 2D datasets in Figure 11. Throughout these datasets, we witness a close resemblance between the final sample distribution and the ground truth target distribution as depicted in Figure 8. The sampling begins with gradual movements across the 2D plane and converges in later diffusion steps. This pattern is further demonstrated by the diffusion value functions, which offer weaker guidance during the initial steps but provide stronger guidance in the final steps. This is in consistency with the exact energy guidance (Lu et al., 2023), which demonstrates that our guidance delivers generation results that exactly conform with the target distribution, while other types of guidance cannot (see Wang et al. (2023) for details).
@@ E.2 Generation Path On Synthetic 2D Tasks
E.3 Training Curves of D4RL Datasets

### v2

@@ Appendix C Proof for the Main Results
Let $p^{\pi_{i}}$ be the diffusion policy at iteration $i$. By Proposition C.4, the sequence $V_{n}^{\pi_{i},s}$ is monotonically increasing. Given that $V^{\pi,s}_{n}$ is finite due to the boundedness of rewards and the pathwise KL, the sequence converges to some $p^{\pi^{*}}$. Since the policy improvement has converged, it must be $V^{\pi^{*},s}_{n}(a)\geq V^{\pi,s}_{n}(a)$ for any diffusion policy $p^{\pi}$, $(s,a)\in\mathcal{S}\times\mathcal{A}$ and $n\in\{0,1,\ldots,N\}$. This completes the proof. ∎
@@ Appendix D Continuous-Time Perspective of the Pathwise KL
Song et al. (2021) build the connection between diffusion models and reverse stochastic differential equations (SDEs) that gradually remove noise from data and transform a prior noise distribution into the target distribution. Based on the SDE understanding of diffusion models, we aim to investigate whether the introduced pathwise KL can be extended to continuous-time diffusion modeling.
We define the forward diffusion SDE, which perturbs data into a prior distribution, as the following:
@@
$D_{\rm KL}\left[\mathbb{P^{\pi}}\|\mathbb{P}^{\nu}\right]$ $=\mathbb{E}_{\mathbb{P}^{\pi}}\left[\log\frac{\mathrm{d}\mathbb{P}^{\pi}}{\mathrm{d}\mathbb{P}^{\nu}}\right]$ (42)
[…]
$=\mathbb{E}_{\mathbb{P}^{\pi}}\left[\int_{0}^{T}\frac{g(t)^{4}\|s^{\pi}(x,t)-s^{\nu}(x,t)\|^{2}}{2g(t)^{2}}\mathrm{d}t\right]$
@@
$D_{\rm KL}\left[\mathbb{P}^{\pi}\|\mathbb{P}^{\nu}\right]$ $=\mathbb{E}_{\mathbb{P}^{\pi}}\left[\int_{0}^{T}\frac{\beta(t)}{2}\|s^{\pi}(x,t)-s^{\nu}(x,t)\|^{2}\mathrm{d}t\right].$ (44)
Eq. (42) presents the analytical form of the KL divergence between two reverse diffusion processes. In the following content, we continue to demonstrate that the pathwise KL introduced in this paper corresponds to Eq. (42) in the limit of $N\to\infty$.
@@
$D_{\rm KL}\left[p^{\pi}_{0:N}\|p^{\nu}_{0:N}\right]$ (45)
$=\mathbb{E}_{p^{\pi}_{0:N}}\left[\sum_{n=1}^{N}D_{\rm KL}\left[p^{\pi,x^{n}}_{n-1|n}\|p^{\nu,x^{n}}_{n-1|n}\right]\right]$
@@
$=\mathbb{E}_{p^{\pi}_{0:N}}\left[\sum_{n=1}^{N}\frac{\beta_{n}(1-\bar{\alpha}_{n})}{2(1-\bar{\alpha}_{n-1})(1-\beta_{n})}\|s_{n}^{\pi}(x^{n})-s_{n}^{\nu}(x^{n})\|^{2}\right],$
@@
$\lim_{N\to\infty}D_{\rm KL}\left[p^{\pi}_{0:N}\|p^{\nu}_{0:N}\right]$ $=\lim_{N\to\infty}\mathbb{E}_{p^{\pi}_{0:N}}\left[\sum_{n=1}^{N}\frac{\beta_{n}(1-\bar{\alpha}_{n})}{2(1-\bar{\alpha}_{n-1})(1-\beta_{n})}\|s_{n}^{\pi}(x^{n})-s_{n}^{\nu}(x^{n})\|^{2}\right]$ (46)
@@
$=\mathbb{E}_{p^{\pi}}\left[\int_{0}^{T}\frac{\beta(t)}{2}\|s^{\pi}(x(t),t)-s^{\nu}(x(t),t)\|^{2}\mathrm{d}t\right],$
@@ E.1 Discussions about Training Time
A systematic breakdown and comparison of the runtime between BDPO, DAC, and Diffusion-QL is illustrated in Figure 11. We evaluate BDPO, DAC, and Diffusion-QL with workstations equipped with NVIDIA RTX 4090 cards and the walker2d-medium-replay-v2 dataset.
@@
E.2 Discussion about Inference Time
Apart from training time, we also evaluate inference time since the latency is critical in real-world applications like robotic control. To benchmark the inference time accurately, we sampled 10,000 states from the walker2d-medium-replay-v2 dataset, executed the diffusion policies with a batch size of 1, and calculated the average inference time across 10,000 trials. Figure 12 compares the inference speed of BDPO and baseline algorithms. Since the specific choice of deep learning backend may have a significant impact on computational latency, we explicitly indicate the backend used for each result.
Among diffusion-based methods, Diffusion-QL generates actions by traversing the reverse diffusion process, and BDPO and DAC generate $N=10$ actions in parallel, followed by Q-value-based sample selection. Given that GPU acceleration effectively amortizes the cost of parallel generation and the calculation of $Q$ values only requires inference through a relatively small network, BDPO and DAC are about 10% slower than JAX-based Diffusion-QL. In the meantime, implementations based on PyTorch are much slower than JAX-based implementations due to the dynamic graph execution. The computational overhead is most pronounced in QGPO, which incurs substantial inference latency due to its classifier-guided generation process.
Figure 12: Inference Time Comparison.
E.3 Generation Path On Synthetic 2D Tasks
We plot the generation results of BDPO on all synthetic 2D datasets in Figure 13. Throughout these datasets, we witness a close resemblance between the final sample distribution and the ground truth target distribution as depicted in Figure 9. The sampling begins with gradual movements across the 2D plane and converges in later diffusion steps. This pattern is further demonstrated by the diffusion value functions, which offer weaker guidance during the initial steps but provide stronger guidance in the final steps. On the contrary, the generation results (Figure 14) of DAC, which leverages $\nabla_{a^{n}}Q^{\pi}(s,a^{n})$ rather than the gradient of diffusion values $\nabla_{a^{n}}V^{\pi,s}_{n}(a^{n})$ to guide the generation, fails to capture the target distribution. This is consistent with the findings from QGPO (Lu et al., 2023), which demonstrates that the guidance in BDPO is exact energy guidance, while the guidance used by DAC is not.
@@ E.3 Generation Path On Synthetic 2D Tasks
Figure 14: Illustration of the diffusion policy and the Q-value function from DAC on synthetic 2D datasets. The regularization strength is set to $\eta=0.06$.
E.4 Training Curves of D4RL Datasets
@@
E.5 Ablation on Policy Parameterization
Full results of different policy parameterizations are presented in Figure 17, from which we found that the diffusion policy achieves the best overall performance. Note that we kept the hyperparameter tuning effort the same across different methods (see Section B.3).
Figure 17: Ablation study with the policy parameterizations. Results are aggregated using 5 independent seeds and 10 evaluation episodes for each seed.
Figure 18: Ablation study with the number of diffusion steps $N$. Results are aggregated using 5 independent seeds and 10 evaluation episodes for each seed.
E.6 Ablation on Diffusion Steps $N$
The number of diffusion steps $N$ impacts both generation quality and divergence calculation. We empirically evaluate three configurations ($N=2$, $5$, and $10$) in Figure 18. Generally, the performance of BDPO degrades substantially when $N$ is too small; however, the improvement diminishes when $N$ exceeds a certain threshold. Based on these observations, we adopt $N=5$ as our default choice to trade off the efficiency and performance.